OpenStudy (anonymous):

how how do i get the standard form of y^2-12y+4x+36=0 by completing the square? plzz help

OpenStudy (anonymous):

hi, this is a parabola

OpenStudy (anonymous):

the standard form is x = a*(y - k)^2 + h

OpenStudy (anonymous):

you complete the square on the y^2 -12y by adding half the linear coefficient squared, that is 36

OpenStudy (anonymous):

this gives you (y-6)^2 +4x =0

OpenStudy (anonymous):

or finally x = (1/4)(y-6)^2

OpenStudy (anonymous):

thank you so much this really helped:)